OpenStudy (anonymous):

please help. log base5 x+logbase5(x+4)=1

OpenStudy (michele_laino):

here we have to apply this property of logarithm: \[\Large {\log _5}x + {\log _5}\left( {x + 4} \right) = {\log _5}\left\{ {x\left( {x + 4} \right)} \right\}\]

OpenStudy (michele_laino):

then, we can rewrite your equation as below: \[\Large {\log _5}\left\{ {x\left( {x + 4} \right)} \right\} = {\log _5}5\] since, by definition of logarithm, we have: \[\Large {\log _5}5 = 1\]

OpenStudy (triciaal):

rephrasing the definition the log of a number is the power to which the base must be raised to get that number